OpenStudy (anonymous):

If g(x) = .5x + 3, find the value of g(0) + 2. I came out with 4, but I'm unsure. Can somebody help me please?

jhonyy9 (jhonyy9):

g(0)=5x0 + 3 = 0 + 3 = 3 g(0) + 2 = 3 + 2 = 5

OpenStudy (saifoo.khan):

oh, i was typing the same thing!

OpenStudy (saifoo.khan):

jhonny is correct.

OpenStudy (anonymous):

thanks

jhonyy9 (jhonyy9):

ok nice ty